# Nonlinear Distortion Radiated from Large Arrays and Active Reconfigurable Intelligent Surfaces

### v1

@@ IV Nonlinear Distortion of an Active RIS
We now consider an active RIS with $N$ reconfigurable elements with the same geometric array structure as in the previous sections. It is deployed to beamform the impinging signal from $K$ single-antenna users to a base station. At the RIS, the time-domain signal received from the users at each reflecting element is phase-shifted and then amplified before being re-radiated. [^3] The order of phase-shifting and amplification has no impact on our results. These amplifiers are nonlinear and this results in a reflected signal that is beamformed towards more locations than there are users, as illustrated in Fig. 2.
@@
Figure 2: The considered setup with an active RIS with $N$ elements that reflects user signals toward a base station while creating nonlinear distortion.
@@ IV-A Phase-shifted Signal
We let the vector $\mathbf{s}_{n}\in\mathbb{C}^{K\times 1}$ contain the transmitted time-domain OFDM symbols from the $K$ users and the matrix $\mathbf{G}\in\mathbb{C}^{M\times K}$ be the channel between the RIS and users. The impinging signal at the RIS will then be $\mathbf{G}\mathbf{s}_{n}$, while the phase-shifted signal $\mathbf{x}_{n}$ at time sample $n$ can be expressed as
$\mathbf{x}_{n}$ $={\bm{\Phi}\mathbf{G}\mathbf{s}_{n}},$ (22)
where the diagonal matrix $\bm{\Phi}$ with unit-modulus entries contains the RIS phase shift configuration. Since the users are assumed to have LoS channels to the RIS, the matrix $\mathbf{G}$ can be written as
$\mathbf{G}=\left[\mathbf{a}\left(\varphi_{1},\theta_{1},r_{1}\right),\dots,\mathbf{a}\left(\varphi_{K},\theta_{K},r_{K}\right)\right],$ (23)
where $\mathbf{a}(\cdot)$ is the array response vector defined in (3) and the RIS observes the $k$-th user from the azimuth angle $\varphi_{k}$, elevation angle $\theta_{k}$ and distance $r_{k}$. We will denote the phase-shift in the $m$-th entry in (3) as
@@
to compensate for the phase-shifts for a signal arriving from the direction $(\varphi_{\tt s},\theta_{\tt s},r_{\tt s})$. The phase-shifted signal in (22) becomes
@@
where $s_{k}[n]$ is the $k$-th entry of the vector $\mathbf{s}_{n}$ and is the time-domain symbol transmitted by the $k$-th user. Hence, the phase-shifted signal at the $m$-th element before amplification is
@@
Figure 3: The PSD of the far-field linear signal (top) and nonlinear distortion (bottom) radiated from an actve RIS with $20\times 20$ antenna elements as a function of the subcarrier index and azimuth angle. All users are seen from the same elevation angle of $\theta_{o}=0^{\circ}$. The vertical lines represent the theoretical azimuth directions of the distortion obtained in Theorem 1.
@@ IV-B Nonlinear Distortion
With the next corollary we provide the spatial characteristics of the third-order nonlinear term (or distortion) radiated from an active RIS.
@@
Suppose the baseband signal in (27) is passed through the non-linear function in (2). Then, the radiation pattern of the third-order non-linear distortion is beamformed in a multitude of directions $(-\varphi_{pqv},-\theta_{pqv},r_{pqv})$ approximated analytically by
$\begin{split}\theta_{pqv}\approx\arcsin\left(\sin\hat{\theta}_{p}-\sin\hat{\theta}_{q}+\sin\hat{\theta}_{v}\right),\end{split}$ (31)
[…]
$\begin{split}r_{pqv}\approx\left(\frac{1}{\hat{r}_{p}}-\frac{1}{\hat{r}_{q}}+\frac{1}{\hat{r}_{v}}\right)^{-1},\end{split}$ (33)
for $p,q,v=1,\ldots,K$, where $\varphi_{pqv}\in[-\frac{\pi}{2},\frac{\pi}{2}]$, $\theta_{pqv}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $r_{pqv}\geq 0$ are the azimuth, elevation and range of each point, respectively.
@@ V Numerical Results
Figure 4: The PSD of the far-field linear signal (top) and nonlinear distortion (bottom) radiated from an active RIS with $20\times 20$ antenna elements as a function of the subcarrier index and elevation angle. All users are seen from the same azimuth angle of $\varphi_{o}=14^{\circ}$. The vertical lines represent the theoretical elevation directions of the distortion obtained in Theorem 1.
(a) Azimuth beamforming
@@
Figure 5: The PSD of the radiative near-field linear signal (top) and nonlinear distortion (bottom) for three different cases of users position. In this example, we assume a UPA $35\times 35$, carrier frequency $f=3$ GHz. In all cases, all users are located at the range of 20m from the ELAA.
In this section, we present simulation results to validate our theoretical analysis and the tightness of the approximations. We consider that both ELAA and RIS are equipped with a UPA with half-wavelength antenna spacing. The parameters of the nonlinear PA model in (2) are chosen such that the error vector magnitude at the output is $3\%$, i.e., $\beta_{1}=1.042$ and $\beta_{3}=-0.0212$. Moreover, without loss of generality, we assume that the RIS phase shifter rotates the impinging signal by $\left(\varphi_{s},\theta_{s}\right)=(-2^{\circ},-4^{\circ})$ to target the base station.
@@ V-B Radiation Pattern
To demonstrate the radiation patterns of the desired (linear) signal and nonlinear distortion, we use the Bussgang decomposition [21] to express the output of the PAs as
$\mathbf{y}_{n}=\mathcal{A}\left(\mathbf{x}_{n}\right)=\mathbf{G}\mathbf{x}_{n}+\mathbf{d}_{n},$ (34)
where the scaled desired signal $\mathbf{u}_{n}\triangleq\mathbf{G}\mathbf{x}_{n}$ is the linear part of the output and the nonlinear distortion is represented by $\mathbf{d}_{n}\in\mathbb{C}^{M\times 1}$, which is uncorrelated with $\mathbf{x}_{n}$. For the considered third-order memoryless PA, the Bussgang gain matrix $\mathbf{G}=\mathbf{I}_{M}+2\beta_{3}\text{diag}(\mathbf{C}_{\bf{yy}}[0])$ is a diagonal matrix, where $\mathbf{C}_{\bf{xx}}[\tau]=\mathbb{E}\{\mathbf{x}_{n}\mathbf{x}^{H}_{n-\tau}\}$ is the covariance matrix of $\mathbf{x}_{n}$. Similarly, the covariance matrix of the distortion vector is given by $\mathbf{C}_{\bf{dd}}(\tau)=\mathbf{C}_{\bf{yy}}[\tau]-\mathbf{C}_{\bf{uu}}[\tau]$, where $\mathbf{C}_{\bf{uu}}[\tau]=\mathbf{G}\mathbf{C}_{\bf{xx}}[\tau]\mathbf{G}$ and the output covariance $\mathbf{C}_{\bf{yy}}[\tau]$ is given analytically in [22, Eq. (27)].
The radiation pattern in the frequency domain is given by the power spectral density (PSD) $\mathbf{S}_{\bf{yy}}(f)$, which is found by taking the Fourier transform of $\mathbf{C}_{\bf{yy}}[\tau]$. Then, the power radiated in the direction of $(\varphi,\theta)$ at frequency $f$ is obtained by multiplying by the corresponding array response vector:
@@ V-C Distortion in the Frequency-Azimuth Domain
(a) Azimuth-depth beamforming
(a) Azimuth-depth beamforming
(b) Finite-depth beamforming
Figure 6: The PSD of the radiative near-field linear signal (top) and nonlinear distortion (bottom) radiated from an ELAA. In this example, we assume a UPA 35 × 35 and carrier frequency f = 3 GHz.
@@ VI Conclusion
In this paper, we analyzed how ELAAs and active RISs with nonlinear amplifiers introduce distortion when transmitting signals. We analyzed the spatial distortion characteristics in the entire 3D space of radiative near-field channels, i.e., in the azimuth, elevation, and depth domains. We derived a new theorem that predicts the beamforming/focusing of the distortion, which match the directions/depths of the desired signals, but distortion also appears at many other distinct locations in the near- and far-field, when the co-scheduled users are located at different near-field distances from the ELAA or RIS. Furthermore, it was shown that when an ELAA transmits or an active RIS receives data signals from multiple users located in different elevation angle directions, the distortion spreads into distinct directions that involve both azimuth and elevation angles not present in the original signal. On the other hand, the azimuth positions of the users do not affect the behavior of distortion in the elevation domain. The nonlinear distortion causes both in-band and out-of-band distortion since the signals expand in the frequency domain.
@@ Appendix A Proof of Theorem 1
In order to find the focal points of the third-order non-linear distortion, we express the radiation pattern of distortion in terms of array factor, i.e.,
${\tt AF}_{3}(\varphi,\theta,r)=\beta_{3}\sum_{m=1}^{M}a_{m}(\varphi,\theta,r)x_{m}[n]|x_{m}[n]|^{2},$ (36)
@@
which is tight when the propagation distance is larger than the $d_{B}=2\Delta$ distance, where $\Delta$ is the maximum linear dimension of the array and the amplitude variations between the point $(\varphi,\theta,r)$ and the aperture of ELAA are negligible[19, 20].
Now, by substituting (27) and (37) into (36), we get the expression
$\begin{split}{\tt AF}_{3}(\varphi,\theta,r)\approx\beta_{3}\sum_{p=1}^{K}\sum_{q=1}^{K}\sum_{v=1}^{K}s_{p}[n]s_{q}^{\ast}[n]s_{v}[n]\\ \times\underbrace{\sum_{m=1}^{M}g_{mpqv}e^{-j\frac{2\pi}{\lambda}\left(\tilde{\phi}_{m}(\varphi,\theta,r)-\tilde{\phi}_{m}(\varphi_{pqv},\theta_{pqv},r_{pqv})\right)}}_{\triangleq{\tt AF}_{pqv}(\varphi,\theta,r)}\end{split}$ (38)
where $\varphi_{pqv}$, $\theta_{pqv}$ and $r_{pqv}$ are given by the expressions (13), (12) and (14), respectively and $g_{mpqv}\triangleq a_{mp}a_{mq}^{\ast}a_{mv}$. Assuming the array response vector in (3) and without loss of generality, we have $g_{mpqv}=1$, $\forall m,p,q,v$. Notice that the amplitude of the term ${\tt AF}_{pqv}(\varphi,\theta,r)$ can be upper bounded using the generalized triangle inequality as
$|{\tt AF}_{pqv}(\varphi,\theta,r)|\leq\sum_{m=1}^{M}\left|e^{-j\frac{2\pi}{\lambda}\left(\tilde{\phi}_{m}(\varphi,\theta,r)-\tilde{\phi}_{m}(\varphi_{pqv},\theta_{pqv},r_{pqv})\right)}\right|,$
where the equality holds if and only if the channel vector equals that of the effective precoder, i.e.,
$\tilde{\phi}_{m}(\varphi,\theta,r)=\tilde{\phi}_{m}(\varphi_{pqv},\theta_{pqv},r_{pqv})$ (39)
Therefore, the term $|{\tt AF}_{pqv}(\varphi,\theta,r)|$ is maximized at the points where $(\varphi,\theta,r)=(\varphi_{pqv},\theta_{pqv},r_{pqv})$ for $p,q,v=1,...,K$.

### v2

@@ IV Nonlinear Distortion of an Active RIS
We now consider an active RIS with $M$ reconfigurable cells with the same geometric array structure as in the previous sections. It is deployed to beamform the impinging signal from $K$ single-antenna users to a base station. At the RIS, the time-domain signal received from the users at each reflecting cell is phase-shifted and then amplified before being re-radiated. [^2] The order of phase-shifting and amplification has no impact on our results. These amplifiers are nonlinear and this results in a reflected signal that is beamformed towards more locations than there are users, as illustrated in Fig. 3.
@@
[…]
Figure 3: The considered setup with an active RIS with $M$ cells that reflects user signals toward a base station while creating nonlinear distortion.
@@ IV-A Phase-shifted Signal
We let the vector $\mathbf{s}_{n}\in\mathbb{C}^{K\times 1}$ contain the transmitted time-domain OFDM symbols from the $K$ users and the matrix $\mathbf{A}\in\mathbb{C}^{M\times K}$ be the channel between the RIS and users as defined in (9). The impinging signal at the RIS will then be $\mathbf{A}\mathbf{s}_{n}$, while the phase-shifted signal $\mathbf{x}_{n}$ at time sample $n$ can be expressed as
$\mathbf{x}_{n}$ $={\bm{\Phi}\mathbf{A}\mathbf{s}_{n}},$ (26)
where the diagonal matrix $\bm{\Phi}$ with unit-modulus entries contains the RIS phase shift configuration. We will denote the phase-shift in the $m$-th entry in (3) as
@@
to compensate for the phase shifts of a signal arriving from user $k$ in the direction $(\varphi_{k},\theta_{k},r_{k})$. The phase-shifted signal in (26) becomes
@@
where $s_{k}[n]$ is the $k$-th entry of the vector $\mathbf{s}_{n}$ and is the time-domain symbol transmitted by the $k$-th user. Hence, the phase-shifted signal at the $m$-th cell before amplification is
@@
Figure 4: The PSD [dBm/Hz] of the far-field linear signal (top) and nonlinear distortion (bottom) radiated from an actve RIS with $20\times 20$ antenna cells as a function of the subcarrier index and azimuth angle. All users are seen from the same elevation angle of $\theta_{o}=0^{\circ}$. The vertical lines represent the theoretical azimuth directions of the distortion obtained in Theorem 1.
@@ IV-B Nonlinear Distortion
With the next corollary, we provide the spatial characteristics of the third-order nonlinear term (or distortion) radiated from an active RIS.
@@
Suppose the baseband signal in (30) is passed through the non-linear function in (2). Then, the radiation pattern of the $p$-th order non-linear distortion is beamformed in a multitude of directions $(-\varphi_{\bm{k}},-\theta_{\bm{k}},r_{\bm{k}})$ approximated analytically by
$\theta_{\bm{k}}$ $\approx\arcsin\left(\sum_{i=0}^{2p}(-1)^{i}\sin\hat{\theta}_{k_{i}}\right),$ (34)
$\varphi_{\bm{k}}$ $\approx\arcsin\left(\frac{1}{\cos\theta_{\bm{k}}}\sum_{i=0}^{2p}(-1)^{i}\sin\hat{\varphi}_{k_{i}}\cos\hat{\theta}_{k_{i}}\right),$ (35)
$r_{\bm{k}}$ $\approx\left(\sum_{i=0}^{2p}(-1)^{i}\frac{1}{\hat{r}_{k_{i}}}\right)^{-1}.$ (36)
where ${\bm{k}}\triangleq\left(k_{0},k_{1},\dots,k_{2p}\right)$ with ${\bm{k}}\in\{1,2,\ldots,K\}^{2p}$, $\varphi_{\bm{k}}\in[-\frac{\pi}{2},\frac{\pi}{2}]$, $\theta_{\bm{k}}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $r_{\bm{k}}\geq 0$ are the azimuth, elevation and range of each focal point, respectively.
@@ V Numerical Results
Figure 5: The PSD [dBm/Hz] of the far-field linear signal (top) and nonlinear distortion (bottom) radiated from an active RIS with $20\times 20$ cells as a function of the subcarrier index and elevation angle. All users are seen from the same azimuth angle of $\varphi_{o}=14^{\circ}$. The vertical lines represent the theoretical elevation directions of the distortion obtained in Theorem 1.
@@
Figure 6: The PSD [dBm/Hz] of the radiative near-field linear signal (top) and nonlinear distortion (bottom) for three different cases of users position. In this example, we assume a UPA $35\times 35$, carrier frequency $f=3$ GHz. In all cases, all users are located at the range of 20m from the ELAA.
In this section, we present simulation results to validate our theoretical analysis and the tightness of the approximations. We consider that both ELAA and RIS are equipped with a UPA with half-wavelength antenna spacing. For simplicity, we assume that the nonlinearities can be modeled by a third-order nonlinear PA model, given by:
$\mathcal{A}(x_{m}[n])=\beta_{1}x_{m}[n]+\beta_{3}x_{m}[n]\lvert x_{m}[n]\rvert^{2},$ (37)
where, unless otherwise specified, the parameters are selected such that the error vector magnitude (EVM) at the output is $3\%$. Specifically, $\beta_{1}=1.042$ and $\beta_{3}=-0.0212$. Moreover, we assume that the RIS phase shifter rotates the impinging signal by $\left(\varphi_{s},\theta_{s}\right)=(-2^{\circ},-4^{\circ})$ to target the base station.
@@ V-B Radiation Pattern
To demonstrate the radiation patterns of the desired (linear) signal and nonlinear distortion, we use the Bussgang decomposition[28] to express the third-order memoryless PA output as
$\mathbf{y}_{n}=\mathbf{G}\mathbf{x}_{n}+\mathbf{d}_{n},$ (38)
where the scaled desired signal $\mathbf{u}_{n}\triangleq\mathbf{G}\mathbf{x}_{n}$ is the linear part of the output while the nonlinear distortion is represented by $\mathbf{d}_{n}\in\mathbb{C}^{M\times 1}$, which is uncorrelated with $\mathbf{x}_{n}$.
The Bussgang gain matrix $\mathbf{G}=\mathbf{I}_{M}+2\beta_{3}\operatorname{diag}(\mathbf{C}_{\mathbf{x}\mathbf{x}}[0])$ is a diagonal matrix, where $\mathbf{C}_{\mathbf{x}\mathbf{x}}[\tau]=\mathbb{E}\{\mathbf{x}_{n}\mathbf{x}_{n}^{\mathrm{H}}\}$ is the covariance matrix of $\mathbf{x}_{n}$. Similarly, the covariance matrix of the distortion vector is given by $\mathbf{C}_{\mathbf{d}\mathbf{d}}(\tau)=\mathbf{C}_{\mathbf{y}\mathbf{y}}[\tau]-\mathbf{C}_{\mathbf{u}\mathbf{u}}[\tau]$, where $\mathbf{C}_{\mathbf{u}\mathbf{u}}[\tau]=\mathbf{G}\mathbf{C}_{\mathbf{x}\mathbf{x}}[\tau]\mathbf{G}$ and the output covariance $\mathbf{C}_{\mathbf{y}\mathbf{y}}[\tau]$ is given analytically in [29, Eq. (27)]. The radiation pattern in the frequency domain is given by the power spectral density (PSD) $\mathbf{S}_{\bf{yy}}(f)$, which is found by taking the Fourier transform of $\mathbf{C}_{\bf{yy}}[\tau]$. Then, the power radiated in the direction of $(\varphi,\theta)$ at frequency $f$ is obtained by multiplying by the corresponding array response vector:
@@
(a) Azimuth-depth beamforming
(b) Finite-depth beamforming
Figure 7: The PSD [dBm/Hz] of the radiative near-field linear signal (top) and nonlinear distortion (bottom) radiated from an ELAA. In this example, we assume a UPA 35 × 35 and carrier frequency f = 3 GHz.
Figure 8: Sum rate as a function of SNR for different levels of EVM with $M=16$ antennas.
Figure 9: Sum rate as a function of SNR for different levels of EVM with $M=100$ antennas.
@@ V-F Distortion in the Depth Domain
V-G Achievable Sum Rates
Next, we examine the impact of nonlinear distortion on the sum rate as a function of the signal-to-noise ratio (SNR), defined as $\text{SNR}\triangleq\rho/\sigma_{n}^{2}$, where $\rho$ represents the received signal power, and $\sigma_{n}^{2}$ denotes the noise power at the receiver. Assuming a LoS channel with no path loss, the received power equals the total transmit power at the BS after amplification, which is considered constant. Under this assumption, we derive the achievable sum rates for MRT and ZF precoding using the following expression:
$R_{\tt sum}=\frac{1}{S}\sum_{k=1}^{K}\sum_{\nu\in S}\log_{2}\left(1+\gamma_{k}[\nu]\right).$ (40)
where $\gamma_{k}[\nu]$ is the signal-to-interference-noise-and-distortion ratio (SINDR), defined as:
$\gamma_{k}[\nu]=\frac{|\mathbf{a}_{k}^{T}\mathbf{G}\mathbf{P}_{\nu,k}|^{2}}{\sum_{i\neq k}|\mathbf{a}_{k}^{T}\mathbf{G}\mathbf{P}_{\nu,i}|^{2}+\mathbf{a}_{k}^{T}\mathbf{S}_{\tt dd}[\nu]\mathbf{a}_{k}^{\ast}+\sigma_{n}^{2}},$ (41)
where $\mathbf{P}_{\nu,k}$ denotes the $k$-th column of the precoder $\mathbf{P}_{\nu}$ (corresponding to the $k$-th user) as defined in (11) and (12). Specifically, Fig. 9 illustrates the sum rate as a function of the SNR for different levels of EVM (i.e., 00 and $3\%$). When the nonlinear distortion is zero (i.e., EVM $=0$), the sum rate increases linearly with the SNR for ZF, as it effectively cancels inter-user interference. However, when introducing an EVM of $3\%$ results in noticeable degradation in the sum rate at high SNR due to the presence of nonlinear distortion, which cannot be mitigated by ZF. Conversely, MRT precoding does not mitigate inter-user interference, which dominates the distortion. Thus, the performance degradation is negligible in the presence of nonlinear distortion. Fig. 9 shows the sum rate as a function of SNR with a system equipped with $M=100$ antennas. Notice that the introduction of nonlinear distortion results in a significant reduction in the sum rate, especially at higher SNR values. With $M=100$ antennas, the performance degradation due to distortion is more pronounced than in the $M=16$ case, demonstrating that the higher the number of antennas, the greater the level of distortion at the UEs. These results clearly demonstrate the impact of EVM on system performance, emphasizing the importance of minimizing EVM to achieve higher sum rates, particularly in scenarios with a larger number of antennas and higher SNR values. The analysis shows that while ZF precoding can handle inter-user interference effectively under ideal conditions, its performance degrades significantly in the presence of nonlinear distortion. On the other hand, MRT precoding shows resilience to nonlinear distortion, but its overall performance is limited by inter-user interference.
Figure 10: Sum rate as a function of SNR for different levels of EVM. In this example, 4 UEs are co-scheduled employing MRT precoding and sub-band scheduling with $M=100$ antennas.
V-H Distortion-Aware Frequency Scheduling
It is possible to use Theorem 1 to predict the beamforming directions of the distortion created by nonlinear amplifiers. This insight can be applied to schedule users in the frequency domain to minimize the influence of distortion. To demonstrate this, we consider a scenario where users are grouped into four clusters, with each cluster containing closely located users, while the clusters themselves are physically far apart. In Fig. 10, 4 users are co-scheduled in a bandwidth of 128 subcarriers, with each user occupying 30 subcarriers. Sub-band scheduling is employed, meaning that different users are allocated distinct parts of the bandwidth.
In the distortion-unaware scheduling method, users are scheduled randomly. This approach often results in co-scheduling users from the same cluster, exposing them to stronger out-of-band (OOB) emissions due to their physical proximity. These emissions exhibit the highest levels of power in the vicinity of the user (cf. Corollary 1), leading to increased distortion and reduced performance (see Figs. 4 and 5).
Conversely, the distortion-aware scheduling method leverages Theorem 1 to predict the beamforming directions of distortion created by nonlinear amplifiers. This enables the minimization of OOB emissions by scheduling users in a way that avoids interference. In this example, by preventing the co-scheduling of users from the same cluster, the method exploits the significant physical separation between clusters to reduce OOB distortion effects that would otherwise be beamformed toward users within the same cluster.
Fig. 10 presents the sum rate as a function of SNR for the two scheduling methods under different EVM levels. The hardware-aware scheduling method achieves a significant sum rate gain compared to the hardware-unaware method, particularly at high SNR values. Moreover, the gain of the hardware-aware method increases with higher non-linear distortion (higher EVM). For instance, at high SNR, the gain is larger for EVM = 10% compared to EVM = 5%, as higher non-linear distortion amplifies the benefits of distortion-aware scheduling. At low SNR values, the performance gap between the two methods is negligible, as non-linear distortion effects are dominated by thermal noise.
@@ VI Conclusion
In this paper, we analyzed how ELAAs and active RISs with nonlinear amplifiers introduce distortion in radiative near-field channels. We examined its spatial characteristics across azimuth, elevation, and depth and derived a theorem predicting its beamforming and focusing behavior. While distortion aligns with the desired signal directions, it also appears at distinct near- and far-field locations when co-scheduled users are at different near-field distances.
We further demonstrated that when multiple users are distributed across different elevation angles, distortion spreads into new azimuth and elevation directions, while user positions in azimuth do not affect distortion in the elevation domain. Additionally, nonlinear amplification causes both in-band and out-of-band spectral expansion, increasing interference in adjacent frequencies.
To mitigate these effects, we proposed distortion-aware scheduling based on our theorem, enabling strategic user allocation in space and frequency to reduce interference. This approach significantly improves system performance, particularly in hardware-impaired scenarios, paving the way for more efficient deployment of ELAAs and active RISs.
@@ Appendix A Proof of Theorem 1
To determine the directions and range of the $p$-th order and the $l$-th memory tap for the non-linear distortion, we express the radiation pattern of the distortion in terms of the array factor, i.e.,
$\begin{split}{\tt AF}_{2p+1}^{(l)}(\varphi,\theta,r)=\beta_{2p+1}[l]\sum_{m=1}^{M}a_{m}(\varphi,\theta,r)\\ \times x_{m}[n-l]|x_{m}[n-l]|^{2p},\end{split}$ (42)
@@
which is tight when the propagation distance is larger than the $d_{B}=2\Delta$ distance, where $\Delta$ is the maximum linear dimension of the array and the amplitude variations between the point $(\varphi,\theta,r)$ and the aperture of ELAA are negligible[26, 27]. Now, by substituting (15) into $x_{m}[n-l]|x_{m}[n-l]|^{2p}$, we get the expression
$\begin{split}x_{m}[n-l]\left|x_{m}[n-l]\right|^{2p}=\left(x_{m}[n-l]\right)^{p+1}\left(x_{m}^{\ast}[n-l]\right)^{p}\\ =\sum_{k_{0}=1}^{K}\sum_{k_{2}=1}^{K}\cdots\sum_{k_{2p}=1}^{K}s_{k_{0}}[n-l]s_{k_{2}}[n-l]\ldots s_{k_{2p}}[n-l]\\ \times e^{-j\frac{2\pi}{\lambda}\sum_{i=0}^{p}\phi_{m}\left(\varphi_{k_{2i}},\theta_{k_{2i}},r_{k_{2i}}\right)}\\ \times\sum_{k_{1}=1}^{K}\sum_{k_{3}=1}^{K}\cdots\sum_{k_{2p-1}=1}^{K}s^{\ast}_{k_{1}}[n-l]s^{\ast}_{k_{3}}[n-l]\ldots s^{\ast}_{k_{2p-1}}[n-l]\\ \times e^{j\frac{2\pi}{\lambda}\sum_{i=1}^{p}\phi_{m}\left(\varphi_{k_{2i-1}},\theta_{k_{2i-1}},r_{k_{2i-1}}\right)}\\ =\sum_{k_{0}=1}^{K}\cdots\sum_{k_{2p}=1}^{K}\underbrace{s_{k_{0}}[n-l]s^{\ast}_{k_{1}}[n-l]\cdots s_{k_{2p}}[n-l]}_{\triangleq s_{{\bm{k}}}[n-l]}\\ \times e^{-j\frac{2\pi}{\lambda}\sum_{i=0}^{2p}(-1)^{i}\phi_{m}\left(\varphi_{k_{i}},\theta_{k_{i}},r_{k_{i}}\right)}.\end{split}$ (44)
Moreover, for notational convenience, we define the phase of the $p$-th order of nonlinear distortion in (44) as
$\tilde{\phi}_{m}(\varphi_{\bm{k}},\theta_{\bm{k}},r_{\bm{k}})\triangleq\sum_{i=0}^{2p}(-1)^{i}\phi_{m}\left(\varphi_{k_{i}},\theta_{k_{i}},r_{k_{i}}\right)$ (45)
and it can be re-written in the form of expression (6) as follows:
$\begin{split}\tilde{\phi}_{m}(\varphi_{\bm{k}},\theta_{\bm{k}},r_{\bm{k}})=k_{z}(m)\underbrace{\sum_{i=0}^{2p}(-1)^{i}\sin\theta_{k_{i}}}_{\triangleq\sin\theta_{\bm{k}}}\\ +k_{y}(m)\underbrace{\sum_{i=0}^{2p}(-1)^{i}\sin\varphi_{k_{i}}\cos\theta_{k_{i}}}_{\triangleq\sin\varphi_{\bm{k}}\cos\theta_{\bm{k}}}\\ -\frac{k^{2}_{z}(m)+k^{2}_{y}(m)}{2}\underbrace{\sum_{i=0}^{2p}(-1)^{i}\frac{1}{r_{k_{i}}}}_{\triangleq\frac{1}{r_{\bm{k}}}}.\end{split}$ (46)
Then by substituting (44) into (42), the array factor of the $p$-th order and the $l$-th memory tap becomes
$\begin{split}{\tt AF}_{2p+1}^{(l)}(\varphi,\theta,r)\approx\beta_{2p+1}[l]\sum_{k_{0}=1}^{K}\cdots\sum_{k_{2p}=1}^{K}s_{\bm{k}}[n-l]\\ \times\underbrace{\sum_{m=1}^{M}e^{-j\frac{2\pi}{\lambda}\left(\tilde{\phi}_{m}(\varphi,\theta,r)-\tilde{\phi}_{m}(\varphi_{\bm{k}},\theta_{\bm{k}},r_{\bm{k}})\right)}}_{\triangleq{\tt AF}_{\bm{k}}(\varphi,\theta,r)},\end{split}$ (47)
where the effective azimuth $\varphi_{\bm{k}}$, elevation $\theta_{\bm{k}}$ and range $r_{\bm{k}}$ of the nonlinear distortion are given Theorem 1. Now, the amplitude of the term ${\tt AF}_{\bm{k}}(\varphi,\theta,r)$ can be upper bounded using the generalized triangle inequality as
$|{\tt AF}_{\bm{k}}(\varphi,\theta,r)|$ $\leq\sum_{m=1}^{M}\left|e^{-j\frac{2\pi}{\lambda}\left(\tilde{\phi}_{m}(\varphi,\theta,r)-\tilde{\phi}_{m}(\varphi_{\bm{k}},\theta_{\bm{k}},r_{\bm{k}})\right)}\right|$
$=M$
Therefore, the equality holds if and only if the channel vector equals that of the effective precoder, i.e.,
$\tilde{\phi}_{m}(\varphi,\theta,r)=\tilde{\phi}_{m}(\varphi_{\bm{k}},\theta_{\bm{k}},r_{\bm{k}})$ (48)
Therefore, the term $|{\tt AF}_{\bm{k}}(\varphi,\theta,r)|$ is maximized at the points where $(\varphi,\theta,r)=(\varphi_{\bm{k}},\theta_{\bm{k}},r_{\bm{k}})$ for $k_{i}=1,2,...,K$ and $i=0,1,...,2p$.